OpenStudy (anonymous):

What is the additive inverse of x + 4 = -9

OpenStudy (anonymous):

We need to isolate the variable, so take off that 4 by subtracting it by both sides.

OpenStudy (anonymous):

Then?

OpenStudy (anonymous):

Well then you get your answer! So subtract 4 :)

OpenStudy (anonymous):

x + 4 = -9 - 4 -4

OpenStudy (anonymous):

I was out for this problem in class thanks

OpenStudy (anonymous):

No problem! Can I haz a medal for helping you :3

OpenStudy (anonymous):

Wait what about 21 + x = 14? Same thing? :o

OpenStudy (anonymous):

Yupp, 21 is positive so the only way to undo something is to do the opposite which is subtract 21 from both sides :)

OpenStudy (anonymous):

Multiplicative inverse? /.\

OpenStudy (anonymous):

I don't know the technical term but I guess yes :/

OpenStudy (anonymous):

No idk that one either, 2/3x=24?

OpenStudy (anonymous):

Well you need to clear the denominator so you would multiply BOTH sides by 3

OpenStudy (anonymous):

Answer would be 72? /.\

OpenStudy (anonymous):

So you get 2x = 72

OpenStudy (anonymous):

No remember you need x by itself

OpenStudy (anonymous):

S how do you undo 2x which is multiplication? (2 * x)

OpenStudy (anonymous):

Opposite of multiplication = ?

OpenStudy (anonymous):

So its 36

OpenStudy (anonymous):

Im sorry, -8 = x/2 im so confused

OpenStudy (anonymous):

Yes! 72 divided by 2 = 36 :D so you get x = 36

OpenStudy (anonymous):

Ok so let me do this, I will only help you when you need it, you will be in charge of doing it, that way you practice, so do you know how to start it?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Remember your primary goal is to get x by itself

OpenStudy (anonymous):

Im so confused

OpenStudy (anonymous):

Ok to get x by itself you need to take out that 2 so how do we do this?

OpenStudy (anonymous):

This is hurting my head idk take it away? Subtract

OpenStudy (anonymous):

X/2 Is division, the only way to undo a division problem is to do the opposite...multiplication! so multiply 2 by both sides to cancel out that pesky 2 :D

OpenStudy (anonymous):

-16=x? /.\

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

You got it :D

OpenStudy (anonymous):

Do you understand any of this or still lost?

OpenStudy (anonymous):

X - (-9) = 12 i know how to do this i just don't know what to do with the -9

OpenStudy (anonymous):

I understand if i see it once

OpenStudy (anonymous):

Ok so - (-9) here we see that 9 is already a negative, so by putting that other - in front of it we can make it 9 because those - cancel out

OpenStudy (anonymous):

So you get X + 9 = 12

OpenStudy (anonymous):

-x = 30? Is it -30?

OpenStudy (anonymous):

Not quite, remember 9 is positive so you have to subtract it from both sides.

OpenStudy (anonymous):

Thats what i did

OpenStudy (anonymous):

12 - 9 = 30?

OpenStudy (anonymous):

This is a completely different problem.-.

OpenStudy (anonymous):

Yes but I am showing you that 12 - 9 is not 30, so you did the problem wrong :)

OpenStudy (anonymous):

Omg! I finished tht problem now i moved on, so im confuse on -x = 30

OpenStudy (anonymous):

OHHHHHHHHHHHHH I get it now haha

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Hmmm... ok I am stumped on this one, I know you need to take out that - sign but I dunno how :/

OpenStudy (ashleyisakitty):

Dubblebarrel please only have one question per thread, and it would be kind to give Meowz a medal for helping you. Thanks all.

OpenStudy (anonymous):

Devise by -x i remember

OpenStudy (anonymous):

How do i do that?

OpenStudy (anonymous):

"Best Response" Button, it's blue, it should be there ---------->

OpenStudy (anonymous):

And what she also means is to close this question and make a new one.

OpenStudy (anonymous):

Its not :o

OpenStudy (anonymous):

I just said its not there

OpenStudy (anonymous):

It doesn't show up for me because it's my own question but it should show up for you by MY post :)

OpenStudy (anonymous):

I just said it is not there

OpenStudy (anonymous):

That's extremely weird, anyways moving on, close this question and open a new one, so you can get more help :)